Which of the following is the equation of a line that passes through (-2, -1) and (-4, -3)?
Answer by stanbon(75887)

Which of the following is the equation of a line
that passes through (-2, -1) and (-4, -3)?
---------
slope = (-3--1)/(-4--2) = -2/-2 = 1
-----
Form: y = mx + b
----
Solve for "b":
-3 = 1*-4+b
b = 1
-----
Equation:
y = m + 1
==================
Cheers,
Stan H.
==================